# On De Giorgi’s lemma for variational interpolants

### v2

@@ 4.3 De Giorgi’s estimate for general dissipation potentials
$(0,\sigma]\ni\rho\mapsto(\widetilde{u}_{\rho},\widetilde{\xi}_{\rho})\in\mathfrak{U}_{{\mathcal{R}}}(\rho,\widetilde{u}_{\rho})\quad\text{such that}$ (4.27)
$\forall\,\rho\in(0,\sigma]\,:\quad{\mathcal{R}}^{*}({-}\widetilde{\xi}_{\rho})=\min\big{\{}\,{\mathcal{R}}^{*}({-}\xi)\>\big{|}\>\xi\in\partial_{\mathcal{R}}{\mathcal{E}}(u^{o},\rho;\widetilde{u}_{\rho})\,\big{\}}=\widehat{\mathscr{C}}_{{\mathcal{R}}}(\rho)\,.$
Obviously, we have
${\mathcal{E}}(\widetilde{u}_{\sigma})+\sigma{\mathcal{R}}\big{(}\frac{1}{\sigma}(\widetilde{u}_{\sigma}{-}u^{o})\big{)}\leq\sup_{u\in\mathscr{U}(\sigma)}\left({\mathcal{E}}(u){+}\sigma{\mathcal{R}}\big{(}\frac{1}{\sigma}(u{-}u^{o})\big{)}\right)\,.$
Thus, from (4.26) we obtain the desired (4.16) for the selection fulfilling (4.27). This concludes the proof of Theorem 4.12.
@@ 4.4 An alternative route to equality via the chain rule
In this section, extending an argument sketched in the introduction we show that, even without radial differentiability for the governing dissipation potential ${\mathcal{R}}$, De Giorgi’s estimate (4.16) can be improved to an identity along a curve $\sigma\mapsto\widetilde{u}_{\sigma}$
@@
Thus, the motivation for this section is to support our conjecture that De Giorgi’s identity is valid in more general situation than those understood by now.
We substantiate the second requirements in the following hypothesis and highlight that, for the variational interpolant it is in fact sufficient to have (piecewise) absolute continuity, whereas the results from Section 4.2 even granted, under additional assumptions, (local) Lipschitz continuity of the curve $\sigma\mapsto\widetilde{u}_{\sigma}$.
Hypothesis 4.14 (Regularity of variational interpolant & chain rule)
Let $\sigma_{*}>0$. We consider a curve $(0,\sigma_{*}]\ni\sigma\mapsto\widetilde{u}_{\sigma}\in J_{\sigma}(u^{o})$ such that
@@
    $\widetilde{u}$ is piecewise absolutely continuous on $(0,\sigma_{*}]$, namely
    $\exists\,p\geq 1\ \exists\,\text{a partition }\{\uptau_{j}\}_{j=0}^{J}\text{ of }[0,\sigma_{*}]\ \forall\,j\in\{1,..,J\}\colon\ \widetilde{u}\in\mathrm{AC}^{p}\big{(}(\uptau_{j-1},\uptau_{j});X\big{)}\,.$ (4.28)
@@
    there exists a measurable selection $\widetilde{\xi}:(0,\sigma_{*}]\to X^{*}$ with $\widetilde{\xi}_{\sigma}\in\partial{\mathcal{E}}(\widetilde{u}_{\sigma})$ for a.a. $\sigma\in(0,\sigma_{*})$ such that
@@
Now, in the upcoming Theorem 4.15 we will consider curves $\sigma\mapsto\widetilde{u}_{\sigma}\in J_{\sigma}(u^{o})$ and $\sigma\mapsto\widetilde{\xi}_{\sigma}\in\partial{\mathcal{E}}(\widetilde{u}_{\sigma})$ as in Hypothesis 4.14 and satisfying De Giorgi’s estimate (4.16). Then, we will clearly have the energy bound $\sup_{\sigma\in]0,\sigma_{*}]}{\mathcal{E}}(\widetilde{u}_{\sigma})<\infty$, as well as
@@
Therefore, if the dual dissipation potential $\xi\mapsto{\mathcal{R}}^{*}(\xi)$ controls $\|\xi\|_{*}^{p^{\prime}}$ with $p^{\prime}$ conjugate to $p$, then combining the above estimate with the condition $u\in\mathrm{AC}^{p}\big{(}(\uptau_{j},\uptau_{j+1});X\big{)}$, we conclude that for the pair $(\tilde{u},\tilde{\xi})$ the second estimate in (4.30) holds.
We are now in a position to state the our last result that De Giorgi’s identity holds even without radial differentiability, if we have some regularity of the variational interpolant $\sigma\mapsto(\widetilde{u}_{\sigma},\widetilde{\xi}_{\sigma})$.
Theorem 4.15 (From De Giorgi’s estimate to De Giorgi’s identity)
Let the $\mathrm{gBGS}$ $(X,{\mathcal{E}},{\mathcal{R}})$ satisfy Hypotheses 3.1. and let $u^{o}\in\mathrm{dom}({\mathcal{E}})$ be fixed. Let $(0,\sigma_{*}]\ni\sigma\mapsto\widetilde{u}_{\sigma}\in J_{\sigma}(u^{o})$ and $(0,\sigma_{*}]\ni\sigma\mapsto\widetilde{\xi}_{\sigma}\in\partial_{\mathcal{R}}{\mathcal{E}}(u^{o},\sigma;\widetilde{u}_{\sigma})$ be as in Hypothesis 4.14. Suppose that they satisfy De Giorgi’s estimate (4.16) on $(0,\sigma_{*}]$.
Then, (4.16) holds as an equality, i.e.
${\mathcal{E}}(\widetilde{u}_{\sigma_{*}})+{\sigma_{*}}{\mathcal{R}}\big{(}\frac{1}{\sigma_{*}}(\widetilde{u}_{\sigma_{*}}{-}u^{o})\big{)}+\int_{0}^{\sigma_{*}}{\mathcal{R}}^{*}(-\widetilde{\xi}_{\rho})\;\!\mathrm{d}\rho={\mathcal{E}}(u^{o})$ (4.31)
@@
$\widetilde{\xi}_{\sigma}\in\mathfrak{A}_{{\mathcal{R}}}(\sigma,\widetilde{u}_{\sigma})=\mathrm{argmin}\big{\{}\,{\mathcal{R}}^{*}({-}\xi)\>\big{|}\>\xi\in\partial_{\mathcal{R}}{\mathcal{E}}(u^{o},\sigma;\widetilde{u}_{\sigma})\,\big{\}}\ \ \text{for a.a.}\,\sigma\in(0,\sigma_{*})\,,$ (4.32)
hence De Giorgi’s identity (3.14) holds.
Proof. Step $1$: Let us fix $j\in\{1,\ldots,J{-}1\}$. It is immediate to check that ${\widetilde{u}_{\uptau_{j}}}^{\pm}\in J_{\uptau_{j}}(u^{o})$, and thus $\Phi_{\uptau_{j}}(u^{o};{\widetilde{u}_{\uptau_{j}}}^{-})=\phi(u^{o};\uptau_{j})=\Phi_{\uptau_{j}}(u^{o};{\widetilde{u}_{\uptau_{j}}}^{+})$, i.e.
${\mathcal{E}}({\widetilde{u}_{\uptau_{j}}}^{-})+\uptau_{j}{\mathcal{R}}\Big{(}\frac{1}{\uptau_{j}}({\widetilde{u}_{\uptau_{j}}}^{-}{-}u^{o})\Big{)}={\mathcal{E}}({\widetilde{u}_{\uptau_{j}}}^{+})+\uptau_{j}{\mathcal{R}}\Big{(}\frac{1}{\uptau_{j}}({\widetilde{u}_{\uptau_{j}}}^{+}{-}u^{o})\Big{)}\,.$ (4.33)
Let now $\sigma\mapsto\widetilde{\xi}_{\sigma}\in\partial_{\mathcal{R}}{\mathcal{E}}(u^{o},\sigma;\widetilde{u}_{\sigma})$ be as in the statement. Hypothesis 4.14 enables us to apply the chain rule on any interval $[s_{*},s^{*}]\subset(\uptau_{j},\uptau_{j+1})$, thus concluding that
@@
$\lim_{\sigma\to\uptau_{j}^{+}}{\mathcal{E}}(\widetilde{u}_{\sigma})=\lim_{\sigma\to\uptau_{j}^{+}}\!\!\left(\phi(u^{o};\sigma){-}\sigma{\mathcal{R}}\Big{(}\frac{1}{\sigma}(\widetilde{u}_{\sigma}{-}u^{o})\Big{)}\right)=\phi(u^{o};\uptau_{j}){-}\uptau_{j}{\mathcal{R}}\Big{(}\frac{1}{\uptau_{j}}(\widetilde{u}_{\uptau_{j}}{-}u^{o})\Big{)}={\mathcal{E}}({\widetilde{u}_{\uptau_{j}}}^{+})$
@@
On the other hand, since $-\widetilde{\xi}_{\sigma}\in\partial{\mathcal{R}}\left(\frac{1}{\sigma}(\widetilde{u}_{\sigma}{-}u^{o})\right)$, by the chain rule for the functional $\sigma\mapsto\sigma{\mathcal{R}}\left(\frac{1}{\sigma}(\widetilde{u}_{\sigma}{-}u^{o})\right)$ we find that
$\uptau_{j+1}{\mathcal{R}}\Big{(}\frac{1}{\uptau_{j+1}}({\widetilde{u}_{\uptau_{j+1}}}^{-}{-}u^{o})\Big{)}-\uptau_{j}{\mathcal{R}}\Big{(}\frac{1}{\uptau_{j}}({\widetilde{u}_{\uptau_{j}}}^{+}{-}u^{o})\Big{)}$ (4.35)
$=\int_{\uptau_{j}}^{\uptau_{j+1}}\Big{(}{\mathcal{R}}\big{(}\frac{1}{\rho}(\widetilde{u}_{\rho}{-}u^{o})\big{)}+\rho\,\big{\langle}{-}\widetilde{\xi}_{\rho},\frac{1}{\rho}\widetilde{u}_{\rho}^{\prime}{-}\frac{1}{\rho^{2}}(\widetilde{u}_{\rho}{-}u^{o})\big{\rangle}\Big{)}\;\!\mathrm{d}\rho\,.$
@@
${\mathcal{E}}({\widetilde{u}_{\uptau_{j+1}}}^{-})+\uptau_{j+1}{\mathcal{R}}\Big{(}\frac{1}{\uptau_{j+1}}({\widetilde{u}_{\uptau_{j+1}}}^{-}{-}u^{o})\Big{)}+\int_{\uptau_{j}}^{\uptau_{j+1}}\!\!\Big{(}\big{\langle}{-}\widetilde{\xi}_{\rho},\tfrac{1}{\rho}(\widetilde{u}_{\rho}{-}u^{o})\big{\rangle}-{\mathcal{R}}\big{(}\frac{1}{\rho}(\widetilde{u}_{\rho}{-}u^{o})\big{)}\Big{)}\;\!\mathrm{d}\rho$
$={\mathcal{E}}({\widetilde{u}_{\uptau_{j}}}^{+})+\uptau_{j}{\mathcal{R}}\Big{(}\frac{1}{\uptau_{j}}({\widetilde{u}_{\uptau_{j}}}^{+}{-}u^{o})\Big{)}\,.$
Now, since $-\widetilde{\xi}_{\rho}\in\partial{\mathcal{R}}\big{(}\frac{1}{\rho}(\widetilde{u}_{\rho}{-}u^{o})\big{)}$, the integrand in the third term on the left-hand side equals ${\mathcal{R}}^{*}({-}\widetilde{\xi}_{\rho})$. In turn, by (4.33), the right-hand side equals $\Phi_{\uptau_{j}}(u^{o};{\widetilde{u}_{\uptau_{j}}}^{-})$. All in all, we conclude that
${\mathcal{E}}({\widetilde{u}_{\uptau_{j+1}}}^{-})+\uptau_{j+1}{\mathcal{R}}\Big{(}\frac{1}{\uptau_{j+1}}({\widetilde{u}_{\uptau_{j}}}^{-}{-}u^{o})\Big{)}+\int_{\uptau_{j}}^{\uptau_{j+1}}{\mathcal{R}}^{*}({-}\widetilde{\xi}_{\rho})\;\!\mathrm{d}\rho$ (4.36)
$\qquad={\mathcal{E}}({\widetilde{u}_{\uptau_{j}}}^{-})+\uptau_{j}{\mathcal{R}}\Big{(}\frac{1}{\uptau_{j}}({\widetilde{u}_{\uptau_{j}}}^{-}{-}u^{o})\Big{)}\quad\text{for all }j\in\{1,\ldots,J{-}1\}\,.$
Step $2$: It remains to derive the analogue of (4.36) for $j=0$. With this aim, we consider the first interval $[0,\uptau_{1}]$ of the partition and fix $\mu\in(0,\uptau_{1}]$. Repeating the arguments from Step $1$ we find
${\mathcal{E}}({\widetilde{u}_{\uptau_{1}}}^{-})+\uptau_{1}{\mathcal{R}}\Big{(}\frac{1}{\uptau_{1}}({\widetilde{u}_{\uptau_{1}}}^{-}{-}u^{o})\Big{)}+\int_{\mu}^{\uptau_{j+1}}{\mathcal{R}}^{*}({-}\widetilde{\xi}_{\rho})\;\!\mathrm{d}\rho$
$={\mathcal{E}}(\widetilde{u}_{\mu})+\mu{\mathcal{R}}\Big{(}\frac{1}{\mu}(\widetilde{u}_{\mu}{-}u^{o})\Big{)}=\Phi_{\mu}(u^{o};\widetilde{u}_{\mu})=\phi(u^{o};\mu)\to{\mathcal{E}}(u^{o})\,,$

### v3

@@ 4.3 De Giorgi estimates for general dissipation potentials
$\sigma{\mathcal{R}}_{\eta}({v}_{\sigma}^{\eta})+E_{0}\leq\sigma{\mathcal{R}}_{\eta}({v}_{\sigma}^{\eta})+{\mathcal{E}}(\widetilde{u}_{\sigma}^{\,\eta})\stackrel{{\scriptstyle(1)}}{{\leq}}{\mathcal{E}}(u^{\circ})\quad\text{for all }\sigma,\,\eta>0\,$, (4.35)
where (1) derives from the minimality of $\widetilde{u}_{\sigma}^{\,\eta}$. Thus, it follows from (4.22) applied with $S:={\mathcal{E}}(u^{\circ})$ that there exists a constant $C_{S}>0$ such that $\sigma\|{v}_{\sigma}^{\eta}\|\leq C_{S}$. Now, recalling that ${v}_{\sigma}^{\eta}=\frac{1}{\sigma}(\widetilde{u}_{\sigma}^{\,\eta}{-}u^{\circ})$ we infer that $\|\widetilde{u}_{\sigma}^{\,\eta}\|\leq\sigma\|{v}_{\sigma}^{\eta}\|\ +\|u^{\circ}\|\leq C_{S}+\|u^{\circ}\|$. All in all, we proved that
$\exists\,\overline{E}>0\ \ \forall\,\eta\in(0,1]\,\ \forall\,\sigma\in(0,\sigma_{*})\,:\qquad{\mathcal{E}}(\widetilde{u}_{\sigma}^{\,\eta})+\|\widetilde{u}_{\sigma}^{\,\eta}\|\leq\overline{E}.\vskip 3.0pt plus 1.0pt minus 1.0pt$ (4.36)
Step 3: Measurable selection. Using the result of Step 2, we are able to apply Proposition 4.15 with the choices $M=B_{\overline{E}}(0)\cap S_{\overline{E}}\subset X$ equipped with the weak topology and ${\mathcal{A}}=\mathcal{S}_{{\mathcal{R}}}$ which is lsc, see (3.11b). Hence, we obtain a measurable function $\sigma\mapsto\widetilde{u}_{\sigma}\in X$ with $\widetilde{u}_{\sigma}^{\,\eta}\rightharpoonup\widetilde{u}_{\sigma}$ along a subsequence $\eta=\eta^{\sigma}_{k}\searrow 0$, where $(\eta^{\sigma}_{k})_{k\in{\mathbb{N}}}$ may depend on $\sigma$. Moreover,
$\mathcal{S}_{{\mathcal{R}}}(\widetilde{u}_{\sigma})\leq\liminf_{\eta\to 0}\mathcal{S}_{{\mathcal{R}}}(\widetilde{u}_{\sigma}^{\,\eta})\quad\text{for all }\sigma\in(0,\sigma_{*}).$ (4.37)
Step 4: Convergence of $\Phi_{\sigma}$. Recall that $\Phi^{\eta}_{\sigma}(u)={\mathcal{E}}(u)+\sigma{\mathcal{R}}_{\eta}\big(\frac{1}{\sigma}(u{-}u^{\circ})\big)$ and that the family $\big(\Phi^{\eta}_{\sigma}(\cdot)\big)_{\eta}$ Mosco-converges to $\Phi_{\sigma}(\cdot)$. Moreover, the functions $\sigma\mapsto\phi^{\eta}(\sigma)=\min\big\{\,\Phi_{\sigma}^{\eta}(u)\>\big|\>u\in X\,\big\}$ are given independently of our curves $\widetilde{u}_{\eta}^{\,\sigma}$ and $\widetilde{u}_{\sigma}$.
From $\widetilde{u}_{\eta}^{\,\sigma}\in\bm{J}_{\eta}(\sigma)$ we have $\phi^{\eta}(\sigma)=\Phi^{\eta}_{\sigma}(\widetilde{u}_{\sigma}^{\,\eta})$. Using the convergence $\widetilde{u}_{\sigma}^{\,\eta^{\sigma}_{k}}\rightharpoonup\widetilde{u}_{\sigma}$ as $k\to\infty$ and the Mosco-convergence of $\Phi^{\eta}_{\sigma}$ we obtain (cf. the argument for (4.30))
$\phi(\sigma)=\Phi_{\sigma}(\widetilde{u}_{\sigma})=\lim_{k\to\infty}\phi^{\eta^{\sigma}_{k}}(\sigma)=\lim_{k\to\infty}\Phi^{\eta^{\sigma}_{k}}_{\sigma}\big(\widetilde{u}_{\sigma}^{\,\eta^{\sigma}_{k}}\big).$
Step 5: The simple De Giorgi estimate. Here we use the trivial lower bound $\mathcal{C}_{{\mathcal{R}}_{\eta}}(\rho;w)\geq{\mathcal{S}}_{{\mathcal{R}}_{\eta}}(w)\geq\mathcal{S}_{{\mathcal{R}}}(w)$ because of ${\mathcal{R}}^{*}_{\eta}\geq{\mathcal{R}}^{*}$. We start from (4.34) and obtain for all fixed $\sigma\in(0,\sigma_{*})$ and all $\eta\in(0,1)$ the estimates
$\Phi_{\sigma}^{\eta}\big(\widetilde{u}_{\eta}^{\,\sigma}\big)+\int_{0}^{\sigma}\mathcal{S}_{{\mathcal{R}}}(\widetilde{u}_{\rho}^{\,\eta})\;\!\mathrm{d}\rho\leq{\mathcal{E}}(u^{\circ}).$
In Step 4 we have shown that the first term converges to the desired limit. For the integral it suffices to use Fatou’s lemma and the liminf estimate (4.37). Thus, we obtain
$\Phi_{\sigma}\big(\widetilde{u}_{\sigma}\big)+\int_{0}^{\sigma}\mathcal{S}_{{\mathcal{R}}}(\widetilde{u}_{\rho})\;\!\mathrm{d}\rho\leq{\mathcal{E}}(u^{\circ}),$
which is the desired simple De Giorgi estimate.
Step 6: The improved De Giorgi estimate. We proceed as in the previous steps, but apply the selection theorem to the enlarged family $(g^{\eta})_{\eta\in(0,1]}$ with
$g^{\eta}:(0,\sigma_{*})\to M_{2}\quad\text{with }\ g^{\eta}_{\sigma}:=(\widetilde{u}_{\sigma}^{\,\eta},\eta)\in M_{2}:=(B_{\overline{E}}(0){\cap}S_{\overline{E}}){\times}[0,1]\subset X{\times}{\mathbb{R}}.$
The measurable function $\mathcal{A}$} is defined via the conditioned ${\mathcal{R}}_{\eta}$ slope as
${\mathcal{A}}\big(\sigma,(u,\eta)\big):=\left\{\begin{array}[]{cl}\mathcal{C}_{{\mathcal{R}}_{\eta}}(\sigma;u)&\text{for }u\in\bm{J}_{\eta}(\sigma),\\ \infty&\text{otherwise},\end{array}\right$.
where for $\eta=0$ we use the notations ${\mathcal{R}}_{0}={\mathcal{R}}$ and $\bm{J}_{0}(\sigma)=\bm{J}(\sigma)$. The additional Hypothesis (4.19) is now exactly tailored in such a way that $\mathcal{A}$} is lower semi-continuous on $(0,\sigma_{*})\to M_{2}$. Hence, Proposition 4.15 provides a selection $\widetilde{u}$ such that
$\widetilde{u}_{\sigma}^{\,\eta}\rightharpoonup\widetilde{u}_{\sigma}\ \text{and }\ \mathcal{C}_{{\mathcal{R}}}(\sigma;\widetilde{u}_{\sigma})\leq\liminf_{\eta\to 0}\mathcal{C}_{{\mathcal{R}}_{\eta}}(\sigma;\widetilde{u}_{\sigma}^{\,\eta})\quad\text{for all }\sigma\in(0,\sigma_{*}).$
As in Step 4 and 5 we can pass to the liminf in (4.34) by combining the last estimate and Fatou’s lemma, and the improved De Giorgi estimate follows.
Finally, the improved estimate with selection 4.21b directly follows via Lemma 4.3.
@@ 4.5 An alternative route to equality via the chain rule
In this section, we extend an argument sketched in the introduction by showing that, even without radial differentiability for the dissipation potential $\mathcal{R}$}, the improved De Giorgi estimate with force selection (4.21b) can be enhanced to an identity along a curve $\sigma\mapsto\widetilde{u}_{\sigma}$
@@
Thus, the motivation for this section is to support our conjecture that the De Giorgi identity is valid in more general situations than those understood by now.
We substantiate the above requirements in the following hypothesis and highlight that, for the variational interpolant it is in fact sufficient to have (piecewise) absolute continuity, whereas the results from Section 4.2 even granted, under additional assumptions, (local) Lipschitz continuity of the curve $\sigma\mapsto\widetilde{u}_{\sigma}$.
Hypothesis 4.16 (Regularity of the variational interpolant & chain rule)
Let $\sigma_{*}>0$. We consider a curve $(0,\sigma_{*})\ni\sigma\mapsto\widetilde{u}_{\sigma}\in J_{\sigma}(u^{\circ})$ such that
@@
    $\widetilde{u}$ is piecewise absolutely continuous on $(0,\sigma_{*})$, namely
    $\hskip-22.76228pt\exists\,p\geq 1\ \exists\,\text{a partition }\{\uptau_{j}\}_{j=0}^{J}\text{ of }[0,\sigma_{*})\ \forall\,j\in\{1,..,J\}\colon\ \widetilde{u}\in\mathrm{AC}^{p}\big((\uptau_{j-1},\uptau_{j});X\big)\,.$ (4.38)
@@
    there exists a measurable selection $\widetilde{\xi}:(0,\sigma_{*})\to X^{*}$ with $\widetilde{\xi}_{\sigma}\in\partial{\mathcal{E}}(\widetilde{u}_{\sigma})$ for a.a. $\sigma\in(0,\sigma_{*})$ such that
@@
Now, in the upcoming Theorem 4.17 we will consider curves $\sigma\mapsto\widetilde{u}_{\sigma}\in J_{\sigma}(u^{\circ})$ and force selections $\sigma\mapsto\widetilde{\xi}_{\sigma}\in\partial_{{\mathcal{R}}}{\mathcal{E}}(\sigma;\widetilde{u}_{\sigma})$ as in Hypothesis 4.16 and satisfying the improved De Giorgi estimate (4.21b) (see Theorem 4.12 for the existence of such selections). Then, we will clearly have the energy bound $\sup_{\sigma\in]0,\sigma_{*})}{\mathcal{E}}(\widetilde{u}_{\sigma})<\infty$, as well as
@@
[…]
We are now in a position to state the our last result that the De Giorgi identity holds even without radial differentiability, if we have some regularity of the variational interpolant $\sigma\mapsto(\widetilde{u}_{\sigma},\widetilde{\xi}_{\sigma})$.
Theorem 4.17 (From the De Giorgi estimate to the identity)
Let the $\mathrm{gBGS}$ $(X,{\mathcal{E}},{\mathcal{R}})$ satisfy Hypotheses 3.1. and let $u^{\circ}\in\mathrm{dom}({\mathcal{E}})$ be fixed. Let $(0,\sigma_{*})\ni\sigma\mapsto\widetilde{u}_{\sigma}\in J_{\sigma}$ and $(0,\sigma_{*})\ni\sigma\mapsto\widetilde{\xi}_{\sigma}\in\partial_{{\mathcal{R}}}{\mathcal{E}}(\sigma;\widetilde{u}_{\sigma})$ satisfy Hypothesis 4.16, as well as the improved De Giorgi estimate with selection (4.21b) on $(0,\sigma_{*})$.
Then, (4.21b) holds as an equality, i.e.
${\mathcal{E}}(\widetilde{u}_{\sigma_{*}})+{\sigma_{*}}{\mathcal{R}}\big(\frac{1}{\sigma_{*}}(\widetilde{u}_{\sigma_{*}}{-}u^{\circ})\big)+\int_{0}^{\sigma_{*}}{\mathcal{R}}^{*}(-\widetilde{\xi}_{\rho})\;\!\mathrm{d}\rho={\mathcal{E}}(u^{\circ})$ (4.41)
@@
$\widetilde{\xi}_{\sigma}\in\mathfrak{A}_{{\mathcal{R}}}(\sigma,\widetilde{u}_{\sigma})=\mathrm{argmin}\big\{\,{\mathcal{R}}^{*}({-}\xi)\>\big|\>\xi\in\partial_{{\mathcal{R}}}{\mathcal{E}}(\sigma;\widetilde{u}_{\sigma})\,\big\}\ \ \text{for a.a.}\,\sigma\in(0,\sigma_{*})\,$, (4.42)
hence the De Giorgi identity (3.17) holds.
Proof. Step 1: Let us fix $j\in\{1,\ldots,J{-}1\}$. It is immediate to check that ${\widetilde{u}_{\uptau_{j}}}^{\pm}\in J_{\uptau_{j}}$, and thus $\Phi_{\uptau_{j}}({\widetilde{u}_{\uptau_{j}}}^{-})=\phi(\uptau_{j})=\Phi_{\uptau_{j}}({\widetilde{u}_{\uptau_{j}}}^{+})$, i.e.
${\mathcal{E}}({\widetilde{u}_{\uptau_{j}}}^{-})+\uptau_{j}{\mathcal{R}}\Big(\frac{1}{\uptau_{j}}({\widetilde{u}_{\uptau_{j}}}^{-}{-}u^{\circ})\Big)={\mathcal{E}}({\widetilde{u}_{\uptau_{j}}}^{+})+\uptau_{j}{\mathcal{R}}\Big(\frac{1}{\uptau_{j}}({\widetilde{u}_{\uptau_{j}}}^{+}{-}u^{\circ})\Big)\,.$ (4.43)
Let now $\sigma\mapsto\widetilde{\xi}_{\sigma}\in\partial_{{\mathcal{R}}}{\mathcal{E}}(\sigma;\widetilde{u}_{\sigma})$ be as in the statement. Hypothesis 4.16 enables us to apply the chain rule on any interval $s_{*},s^{*}]\subset(\uptau_{j},\uptau_{j+1})$, thus concluding that
@@
$\lim_{\sigma\to\uptau_{j}^{+}}{\mathcal{E}}(\widetilde{u}_{\sigma})=\lim_{\sigma\to\uptau_{j}^{+}}\!\!\left(\phi(\sigma){-}\sigma{\mathcal{R}}\Big(\frac{1}{\sigma}(\widetilde{u}_{\sigma}{-}u^{\circ})\Big)\right)=\phi(\uptau_{j}){-}\uptau_{j}{\mathcal{R}}\Big(\frac{1}{\uptau_{j}}(\widetilde{u}_{\uptau_{j}}{-}u^{\circ})\Big)={\mathcal{E}}({\widetilde{u}_{\uptau_{j}}}^{+})$
@@
In turn, since $\widetilde{\xi}_{\sigma}\in\partial_{{\mathcal{R}}}{\mathcal{E}}(\sigma;\widetilde{u}_{\sigma})$ implies $\widetilde{\xi}_{\sigma}\in\partial{\mathcal{R}}\left(\frac{1}{\sigma}(\widetilde{u}_{\sigma}{-}u^{\circ})\right)$, we may apply the chain rule for the functional $\sigma\mapsto\sigma{\mathcal{R}}\left(\frac{1}{\sigma}(\widetilde{u}_{\sigma}{-}u^{\circ})\right)$ and find that
$\uptau_{j+1}{\mathcal{R}}\Big(\frac{1}{\uptau_{j+1}}({\widetilde{u}_{\uptau_{j+1}}}^{-}{-}u^{\circ})\Big)-\uptau_{j}{\mathcal{R}}\Big(\frac{1}{\uptau_{j}}({\widetilde{u}_{\uptau_{j}}}^{+}{-}u^{\circ})\Big)$ (4.45)
$=\int_{\uptau_{j}}^{\uptau_{j+1}}\Big({\mathcal{R}}\big(\frac{1}{\rho}(\widetilde{u}_{\rho}{-}u^{\circ})\big)+\rho\,\big\langle{-}\widetilde{\xi}_{\rho},\frac{1}{\rho}\widetilde{u}_{\rho}^{\prime}{-}\frac{1}{\rho^{2}}(\widetilde{u}_{\rho}{-}u^{\circ})\big\rangle\Big)\;\!\mathrm{d}\rho\,.$
@@
${\mathcal{E}}({\widetilde{u}_{\uptau_{j+1}}}^{-})+\uptau_{j+1}{\mathcal{R}}\Big(\frac{1}{\uptau_{j+1}}({\widetilde{u}_{\uptau_{j+1}}}^{-}{-}u^{\circ})\Big)+\int_{\uptau_{j}}^{\uptau_{j+1}}\!\!\Big(\big\langle{-}\widetilde{\xi}_{\rho},\tfrac{1}{\rho}(\widetilde{u}_{\rho}{-}u^{\circ})\big\rangle-{\mathcal{R}}\big(\frac{1}{\rho}(\widetilde{u}_{\rho}{-}u^{\circ})\big)\Big)\;\!\mathrm{d}\rho$
$={\mathcal{E}}({\widetilde{u}_{\uptau_{j}}}^{+})+\uptau_{j}{\mathcal{R}}\Big(\frac{1}{\uptau_{j}}({\widetilde{u}_{\uptau_{j}}}^{+}{-}u^{\circ})\Big)\,.$
Now, since $\widetilde{\xi}_{\rho}\in\partial{\mathcal{R}}\big(\frac{1}{\rho}(\widetilde{u}_{\rho}{-}u^{\circ})\big)$, the integrand in the third term on the left-hand side equals ${\mathcal{R}}^{*}({-}\widetilde{\xi}_{\rho})$. In turn, by (4.43), the right-hand side equals $\Phi_{\uptau_{j}}({\widetilde{u}_{\uptau_{j}}}^{-})$. All in all, we conclude that
${\mathcal{E}}({\widetilde{u}_{\uptau_{j+1}}}^{-})+\uptau_{j+1}{\mathcal{R}}\Big(\frac{1}{\uptau_{j+1}}({\widetilde{u}_{\uptau_{j}}}^{-}{-}u^{\circ})\Big)+\int_{\uptau_{j}}^{\uptau_{j+1}}{\mathcal{R}}^{*}({-}\widetilde{\xi}_{\rho})\;\!\mathrm{d}\rho$ (4.46)
$\qquad={\mathcal{E}}({\widetilde{u}_{\uptau_{j}}}^{-})+\uptau_{j}{\mathcal{R}}\Big(\frac{1}{\uptau_{j}}({\widetilde{u}_{\uptau_{j}}}^{-}{-}u^{\circ})\Big)\quad\text{for all }j\in\{1,\ldots,J{-}1\}\,.$
Step 2: It remains to derive the analogue of (4.46) for $j=0$. With this aim, we consider the first interval $0,\uptau_{1}]$ of the partition and fix $\mu\in(0,\uptau_{1}]$. Repeating the arguments from Step 1 we find
${\mathcal{E}}({\widetilde{u}_{\uptau_{1}}}^{-})+\uptau_{1}{\mathcal{R}}\Big(\frac{1}{\uptau_{1}}({\widetilde{u}_{\uptau_{1}}}^{-}{-}u^{\circ})\Big)+\int_{\mu}^{\uptau_{j+1}}{\mathcal{R}}^{*}({-}\widetilde{\xi}_{\rho})\;\!\mathrm{d}\rho$
$={\mathcal{E}}(\widetilde{u}_{\mu})+\mu{\mathcal{R}}\Big(\frac{1}{\mu}(\widetilde{u}_{\mu}{-}u^{\circ})\Big)=\Phi_{\mu}(\widetilde{u}_{\mu})=\phi(\mu)\to{\mathcal{E}}(u^{\circ})\,$,